Curriculum Vitae

William Gordon McCallum

Address

Department of Mathematics University of Arizona Tucson, AZ 85721, U.S.A. Phone: 520-440-1729 Email: wmc@math.arizona.edu Homepage: www.math.arizona.edu/˜ wmc

Education

1985 Ph. D., Harvard University, under the direction of Barry Mazur

1977 B. Sc., University of New South Wales, Sydney

Employment (last 20 years)

2016– President, Illustrative Mathematics

2014–2016 President and CEO, Illustrative Mathematics

2009–13 Head, Department of Mathematics

2006–9 Director, Institute for Mathematics and Education

2006– University Distinguished Professor, University of Arizona (on leave since 2013)

1996–2005 Professor, University of Arizona

Visiting Positions

2006/1–5 Fellow of the Udall Center for Studies in Public Policy, University of Arizona

2001/3 Prof. Invité, Université Joseph Fourier, Grenoble 2001/2 Prof. Invité, Université de Paris-Sud, Orsay 2000/9–12 Visitor, Institut des Hautes Études Scientifiques, 1996/7 Visitor, Institut des Hautes Études Scientifiques 1995–96 Fellow, Institute for Advanced Study 1995/7 Prof. Invité, Université de Bordeaux I 1993–94 Visitor, Institut des Hautes Études Scientifiques 1991/1–6 Visiting Lecturer, Macquarie University, Sydney

Honors and Awards

2012 Mary P. Dolciani Award, Mathematical Association of America 2012 AMS Award for Distinguished Public Service 2006 University of Arizona College of Science Galileo Circle Fellow 2005 NSF Director’s Award for Distinguished Teaching Scholars 1996 University of Arizona College of Science Innovation in Teaching

Award

1995 American Mathematical Society Centennial Research Fellowship

Published articles

1. (with Bjorn Poonen) The method of Coleman and Chabauty. In K. Belabas, editor, Explicit Methods in Number Theory: Rational Points and Diophantine Equations, Soc. Math. de France, 2013.

2. Restoring and Balancing In Usiskin, Anderson, and Zotto (eds), Future Curricular Trends in School Algebra and Geometry, Information Age Publishing (2010)

3. (with Benjamin Levitt) Yet more elements in the Shafarevich-Tate group of the jacobian of a Fermat curve. In Lauter, Kristin E. and Ribet, Kenneth A. (eds), Computational Algebraic Geometry, Contemporary Mathematics, Vol. 463. AMS, 2008.

4. Assessing the strands of student proficiency in elementary algebra. In Alan H. Schoenfeld, editor, Assessing Mathematical Proficiency, volume 53 of Mathematical Sciences Research Institute Publications, New York, 2007. Cambridge University Press.

5. Computer algebra and human algebra. In Ki Hyoung Ko and Dean Arganbright, editors, Enhancing University Mathematics, volume 14 of CBMS Issues in Mathematics Education, pages 43–50, Rhode Island, 2007. American Mathematical Society.

6. Promoting work on education in mathematics departments. Notices of the AMS, pages 1093–1098, October 2003.

7. (with Pavlos Tzermias) On Shafarevich-Tate groups and the arithmetic of Fermat curves. In Number theory and algebraic geometry, volume 303 of London Math. Soc. Lecture Note Ser., pages 203–226. Cambridge Univ. Press, Cambridge, 2003.

8. (with Romyar T. Sharifi) A cup product in the Galois cohomology of number fields. Duke Math. J., 120(2):269–310, 2003.

9. (with Randall M. Richardson) The third R in literacy. In Bernard L. Madison and Lynn Arthur Steen, editors, Quantitative Literacy: Why Numeracy Matters for Schools and Colleges, pages 99–106, Princeton, 2003. NCED.

10. Thinking out of the box. In James T. Fey, editor, Computer Algebra Systems in Secondary School Mathematics Education. NCTM, 2003.

11. Greenberg’s conjecture and units in multiple ℤ_{p}-extensions. Amer. J. Math., 123:909–930, 2001.

12. Brauer points on fermat curves. Bull. Austral. Math. Soc., 63:393–406, 2001.

13. (with Sang Yook An, Seog Young Kim, David C. Marshall, Susan H. Marshall, and Alexander R. Perlis) Jacobians of genus one curves. J. Number Theory, 91(2):304–315, 2001.

14. The goals of the calculus course. In Susan L. Ganter, editor, Calculus Renewal, Issues for Undergraduate Education in the Next Decade, chapter 2, pages 11–22. Kluwer Academic, New York, 2000.

15. Will this be on the exam? In Steven G. Krantz, editor, How to Teach Mathematics, pages 233–240. American Mathematical Society, Providence, 1999.

16. A duality theorem in the multivariable Iwasawa theory of local fields. J. reine angew. Math., 464:143–172, 1995.

17. On the method of Coleman and Chabauty. Math. Ann., 299:565–596, 1994.

18. The arithmetic of Fermat curves. Math. Ann., 294:503–511, 1992.

19. Rigor in the undergraduate calculus curriculum. Notices of the AMS, 38(9):1131–1132, 1991.

20. Kolyvagin’s work on Shafarevich-Tate groups. In J. Coates and M. J. Taylor, editors, L-functions and Arithmetic, Proceedings of the Durham Symposium, July, 1989, volume 153 of LMS Lecture Note Series, pages 295–316, Cambridge, 1991. Cambridge University Press.

21. Tate duality and wild ramification. Math. Ann., 288:553–558, 1990.

22. On the Shafarevich-Tate group of the Jacobian of a quotient of the Fermat curve. Invent. Math., 93(3):637–666, 1988.

23. (with Robert Coleman) Stable reduction of Fermat curves and Jacobi sum Hecke characters. J. reine angew. Math., pages 41–101, 1988.

24. Duality theorems for Nèron models. Duke Math. Journal, 53:1093–1124, 1986.

25. The degenerate fibre of the Fermat curve. In Number theory related to Fermat’s last theorem (Cambridge, Mass., 1981), volume 26 of Progr. Math., pages 57–70. Birkhäuser Boston, Mass., 1982.

Books

1. Algebra: Form and Function, McCallum, Connally, Hughes-Hallett et al., John Wiley & Sons, New York, 2010.

2. Applied Calculus, Fraser-Locke, Hughes-Hallett, Gleason et al., John Wiley & Sons, New York, four editions, 1999–.

3. Multivariable Calculus, McCallum, Hughes-Hallett, Gleason et al., John Wiley & Sons, New York, four editions, 1997–

4. Calculus, Hughes-Hallett, Gleason et al., John Wiley & Sons, New York, five editions, 1993–.

5. Teaching Mathematics in Colleges and Universities: Case Studies for Today’s Classroom, by Solomon Friedberg. Contributor with many others, American Mathematical Society, 2001.

6. Contemporary Issues in Mathematics Education, proceedings of a conference held at MSRI in December 1996, co-editor with Steven Krantz and Estela Gavosto, Cambridge University Press, 1999

Grants since 2010

2014–15 Grant for website redevelopment, the Helmsley Charitable Trust.

2014–15 Organizational Effectiveness Grant, The William and Flora Hewlett Foundation.

2013– Middle School Blueprints, 100%, The William and Flora Hewlett

Foundation.

2013–2014 High School Blueprints, 100%, The William and Flora Hewlett

Foundation.

2011–15 The Illustrative Mathematics Project, 100%, The Bill and Melinda

Gates Foundation

2010– Progressions Documents for the Common Core Math Standards, 100%, Brookhill Foundation

Selected Service

2013–2018 AMS Committee on the Profession 2013–2019 MAA Committee on the James R. C. Leitzel Lecture. 2008–2013 MAA Committee on the Mathematical Education of Teachers. 2009–10 Chair of work team and lead writer for the Common Core State

Standards in Mathematics

2008– Cochair of update committee for Achieve high school benchmarks 2008–10 Chair of Conference Board of the Mathematical Sciences 2007–8 Revisions committe for Arizona state standards 2007–8 Writing group for NCTM high school curriculum document